I have to solve for (u) but I forgot how to do this problem...
-1/2=3/7u+1/3

Answer by rfer(16322)

-1/2=3/7u+1/3
-1/2-1/3=3/7u
-3/6-2/6=3/7u
-5/6=3/7u
u=-5/6*7/3
u=-35/42
u=5/6